# The Born rule – 100 years ago and today

### v2

@@
Acknowledgment. I recall with pleasure heated discussions with Francois Ziegler, several years ago, about the detailed history of the Born rule.
@@ 1 Ten formulations of the Born rule
For example, we read in Jeans [35, p.91] (1924): ”Obviously quantum-dynamics must be intimately concerned with the probability of such jumps. […] measure, in some way at present unknown, the probabilities of jumps in the velocity and perhaps also in the position of an electron which forms part of an atomic system.” And in Bohr et al. [7, p.75] (1924), we read: ”Ferner nehmen wir an, dass das Vorkommen von Übergangsprozessen, sowohl für das gegebene Atom selbst wie für die anderen Atome, mit denen es kommuniziert, mit diesem Mechanismus durch Wahrscheinlichkeitsgesetze verknüpft ist, die den Gesetzen der Einsteinschen Theorie für die von äusserer Strahlung induzierten Übergänge zwischen stationären Zuständen analog sind.”
In February 1926, this interpretation was already semiquantitative, valid for large quantum numbers: ”Nach der Quantentheorie kann sich der innere Zustand eines Atoms nur sprunghaft ändern. Es sind nur die diskreten stationären Zustände der Atome möglich. An Stelle der klassischen kontinuierlichen Energieaufnahme haben wir daher eine diskontinuierliche in einzelnen Sprüngen, für die gewisse Übergangswahrscheinlichkeiten bestehen. Es müssen dann zwischen beiden korrespondenzmässige Beziehungen vorhanden sein, d.h. näherungsweise, für grosse Quantenzahlen sogar exakt, Übereinstimmung zwischen der klassischen und quantentheoretischen Rechnung herrschen.” (Nordheim [60, p.512])
@@
In June 1926, Born [9, p.822] found a highly successful statistical interpretation of the then very new wave function of Schrödinger, from which he could derive very specific physical conclusions, namely the qualitative features of collision experiments (which were later verified quantitatively). This earned Born the Nobel prize in 1954 ”for his fundamental research in quantum mechanics, especially for his statistical interpretation of the wavefunction” [59].
Wallner’s Nobel prize presentation speech for the general public comments ([59]): ”Schrödinger had not solved the problem of how it is possible to make statements about the positions and velocities of particles if one knows the wave corresponding to the particle. Born provided the solution to the problem. He found that the waves determine the probability of the measuring results. For this reason, according to Born, quantum mechanics gives only a statistical description.” This is a bit inaccurate, as Born didn’t find probabilities for positions and velocities of particles (this was done a little later by Pauli) but probabilities for state transitions of particles. This motivates a detailed study of the history of the statistical interpretation of quantum mechanics.
@@ 1.1 The Born rule before 1970
The Born rule has many slightly inequivalent formulations, ten different forms of which we state in the following, in modern terminology. The universal form (BR-US) and the discrete form (BR-DS) of the Born rule are taken almost verbatim from Wikipedia [74]. Four other formulations. are taken with minor changes from Chapter 14 of Neumaier [52] (but with the scattering form restated in two almost identical formulations), and one more from Part II of Neumaier & Westra [54]. Two more general formulations were discovered only after 1970, and are described in Subsection 1.2. Finally, a tenth form (BR-Q) is presented in Subsection 1.3, adapted to quantum statistical mechanics and quantum field theory.
(BR-OSc) Born rule (objective scattering form): In a scattering experiment described by the S-matrix $S$,
@@
(BR-OE) Born rule (objective expectation form): The value of a quantity corresponding to an operator $X$ of a system in the pure state $\psi$ or the mixed state $\rho$ equals on average the quantum expectation value
@@
respctively.
(BR-MSc) Born rule (measured scattering form): In a scattering experiment described by the S-matrix $S$, formula (1) gives the conditional probability density that when particles prepared in the in-state $\psi_{{{\mbox{\rm\scriptsize in}}}}$ are scattered, they are found in the out-state $\psi_{{{\mbox{\rm\scriptsize out}}}}$. Here the in- and out-states are asymptotic eigenstates of the Hamiltonian, labelled by a maximal collection of independent quantum numbers (including the energies, momenta and angular momenta of the particles involved).
@@
(BR-US) Born rule (universal spectral form): If a quantity corresponding to a self-adjoint Hermitian operator $X$ is measured in a system described by a pure state with normalized wave function $\psi$ then
@@
Since the formulation involves projection operators, measurements satisfying (BR-US) are called projective measurements; another name for these is von Neumann measurements. Not all measurements are of this kind; see Subsection 1.2 for a more general class of measurements, and Subsection 3.4 for a fuller discussion of the practical limitations the class of projective measurements has.
(BR-FS) Born rule (finite spectral form): In a projective measurement of a Hermitian, self-adjoint operator $X$ (or vector of pairwise commuting operators) with a finite spectrum, the possible values are precisely the finitely many eigenvalues $\lambda_{k}$ of $X$ (or joint eigenvalues of the components of $X$), measured with a probability of
@@
(BR-DS) Born rule (discrete spectral form): Suppose that a quantity corresponding to a self-adjoint Hermitian operator $X$ with a discrete spectrum is measured in a system described by a pure state with normalized wave function $\psi$. Then:
@@
(BR-ME) Born rule (measured expectation form): If a quantity corresponding to a self-adjoint Hermitian operator $X$ is measured on a system in the pure state $\psi$ or the mixed state $\rho$, the results equal on average the quantum expectation value (2) or (3), respectively.
@@ 1.2 The Born rule after 1970
This straightforward generalization of the concept of a finite probability measure, given by a family of finitely many nonnegative numbers $p_{k}$ (probabilities) summing to 1, is a simplified version of a finite POVM. The following generalization of the finite spectral form (BR-FS) of the Born rule is essentially in Ali & Emch [1, (2.14)]:
(BR-POVM) Born rule (POVM form): In a general quantum measurement, the possible values are finitely many distinct numbers or vectors $\lambda_{k}$, measured with a probability of (4), where the $P_{k}$ form a quantum measure.
@@
To justify (BR-POVM), the usual practice is to assume the Born rule for projective measurements as a basic premise, with a purely historical justification. Later (if at all), the more general quantum measure (POVM) setting is postulated and justified (e.g., in Fuchs & Peres [30]) in terms of the Born rule for projective measurements, not in the physical state space of the system plus its environment, but in an unphysical state space formally constructed using an appropriate ancilla on the basis of Naimark’s theorem (Naimark [49]). This shows the consistency with the Born rule, but the derivation is physically artificial since it is not solely based on states in the physical Hilbert space.
Following Neumaier & Westra [54, Part II], we give in Section 3 precise definitions that lead to the correspondence between statistical expectations and quantum expectation claimed by the following condensed form of the Born rule.
(BR-C) Born rule (condensed form): The statistical expectation of the measurement results equals the quantum expectation of the measured quantity.
This form of the Born rule is derived in Subsection 3.3 together with (BR-POVM) from a new, nontechnical measurement principle that is easy to motivate and understand, without any technicalities regarding the spectrum of operators.
@@ 1.3 Quantum expectation values
In quantum field theory, quantum expectation values of non-Hermitian operators are routinely used; they are already needed in the definition of so-called $N$-point functions as quantum expectation values of products of field operators. That $2$-point functions are in principle observables through linear response theory shows that the operators corresponding to observable quantities need not be Hermitian. Indeed, Dirac’s 1930 quantum mechanics textbook (Dirac [21, p.28]), which introduced the name ”observable” for operators, used this terminology for arbitrary linear operators. It is therefore interesting to note that all later editions of Dirac’s textbook and all later textbooks on quantum mechanics require the restriction to Hermitian operators possessing a real spectral resolution, i.e., in modern terminology, to self-adjoint Hermitian operators. Probably Dirac became aware of the fact that when $X$ is not normal (and in particular when $X$ is defective, hence has not even a spectral resolution), it is impossible to give the quantum value $\langle X\rangle$ of $X$ a spectral interpretation – which before 1970 was an essential ingredient of all measurement-based forms of the Born rule. Only the more recent formulations (BR-POVM), (BR-C), and (BR-Q) are again free of a spectral restriction.
In quantum statistical mechanics and quantum field theory one usually makes a theoretical analysis of quantum processes where no actual measurement is made before the process is completed. Since one cannot invoke formulations of the Born rule that make reference to measurement, all probabilities and expectation values that appear in such an analysis, must be considered as formal expressions (2)–(4), with a clear mathematical meaning independent of the interpretation of quantum mechanics. However, since the objective formulations of the Born rule are inconsistent, these quantum probabilities and quantum expectation values are purely theoretical quantities without any direct relation to reality.
If one nevertheless wants to give the computations in quantum statistical mechanics and quantum field theory an objective physical meaning, one must therefore define the objective quantum value of a vector $X$ of operators to be the quantum expectation value $\langle X\rangle$:
(BR-Q) Born rule (quantum value form): The quantum value $\overline{X}$ of a quantity corresponding to an operator $X$ of a system in the pure state $\psi$ or the mixed state $\rho$ is defined as the quantum expectation value (2) or (3), respectively. It is approximately measured, with an uncertainty of
@@
This form of the Born rule expresses the spirit of Heisenberg’s statement (in his 1927 paper on the uncertainty relation): ” Jeder quantentheoretischen Grösse oder Matrix lässt sich eine Zahl, die ihren ’Wert’ angibt, mit einem bestimmten wahrscheinlichen Fehler zuordnen; der wahrscheinliche Fehler hängt vom Koordinatensystem ab; für jede quantentheoretische Grösse gibt es je ein Koordinatensystem, in dem der wahrscheinliche Fehler für diese Grösse verschwindet. (Heisenberg [31, p.181f])
@@ 1.4 Discussion
Since (BR-OSc) and (BR-OE) do not refer to measurement, they are not loaded with the difficulties surrounding the measurement problem of quantum mechanics. However, as we shall discuss in more detail in Subsection 2.3, they are mathematically inconsistent and hence were abandoned already by 1930.
@@
tacitly made in classical statistical physics, and imported from there into the spectral forms of the Born rule. In practice, we only measure a small set of distinguished variables – distances, position, energy, momentum, angular momentum, spin, helicity, and not arbitrary nonlinear expressions in these. Hence it is quite reasonable to require (D) not for nonlinear $f$ but only for linear $f$. Indeed, (D) is also incompatible with (BR-POVM). Giving up (D) is necessary if one wants to account for experiments involving the simultaneous measurement of position and momentum.
The three newest forms of the Born rule, (BR-POVM), (BR-C) , and (BR-Q), are mathematically much more elementary since they do not depend on the notion of self-adjointness of operators. As already stated in Subsection 1.2, the POVM form (BR-POVM) of the Born rule is a generalization of the finite spectral form (BR-FS). It is more or less equivalent to the condensed form (BR-C); see Subsection 3.3.
@@
von Neumann [56, p.255] derives for a quantum expectation value satisfying four (for von Neumann) plausible conditions (conditions A.–D. listed and motivated on p.250–252) the necessity of the formula (3) with a Hermitian density operator $\rho$ (his $U$) of trace 1. This is abstract mathematical reasoning, still within the core of formal (uninterpreted) quantum mechanics, independent of any relation to measurement. Note that assuming von Neumann’s condition D. (which is a restricted form of condition (D) above, forces the measurements to be projective measurements. Thus he misses the POVM generalization of the Born rule since the latter violates his condition D.

### v3

@@
Acknowledgment. I recall with pleasure heated discussions with Francois Ziegler, several years ago, about the detailed history of the Born rule. I also want to thank Roger Balian for discussions related to this paper, especially about the condensed form (BR-C) of Born’s rule and its use in the maximum entropy principle.
@@ 1 Ten formulations of the Born rule
For example, we read in Jeans [39, p.91] (1924): ”Obviously quantum-dynamics must be intimately concerned with the probability of such jumps. […] measure, in some way at present unknown, the probabilities of jumps in the velocity and perhaps also in the position of an electron which forms part of an atomic system.” And in Bohr et al. [8, p.75] (1924), we read: ”In addition, we assume that the occurrence of transition processes, both for the given atom itself and for the other atoms, with which it communicates, is linked with this mechanism through probability laws that are analogous to the laws of Einstein’s theory for the transitions induced by external radiation between stationary states.”
(German original: ”Ferner nehmen wir an, dass das Vorkommen von Übergangsprozessen, sowohl für das gegebene Atom selbst wie für die anderen Atome, mit denen es kommuniziert, mit diesem Mechanismus durch Wahrscheinlichkeitsgesetze verknüpft ist, die den Gesetzen der Einsteinschen Theorie für die von äusserer Strahlung induzierten Übergänge zwischen stationären Zuständen analog sind.”)
In February 1926, this interpretation was already semiquantitative, valid for large quantum numbers: ”According to the quantum theory, the inner state of an atom can change only in sudden jumps. Only the discrete stationary states of the atoms are possible. In place of the classical continuous absorption of energy we therefore have a discontinuous one in single jumps, for which certain transition probabilities exist. Between both must then exist correspondence relations, i.e., an approximate, for large quantum numbers even exact, agreement between the classical and the quantum theoretical calculation.”
(German original, Nordheim [64, p.512]: ”Nach der Quantentheorie kann sich der innere Zustand eines Atoms nur sprunghaft ändern. Es sind nur die diskreten stationären Zustände der Atome möglich. An Stelle der klassischen kontinuierlichen Energieaufnahme haben wir daher eine diskontinuierliche in einzelnen Sprüngen, für die gewisse Übergangswahrscheinlichkeiten bestehen. Es müssen dann zwischen beiden korrespondenzmässige Beziehungen vorhanden sein, d.h. näherungsweise, für grosse Quantenzahlen sogar exakt, Übereinstimmung zwischen der klassischen und quantentheoretischen Rechnung herrschen.”)
@@
In June 1926, Born [10, p.822] found a highly successful statistical interpretation of the then very new wave function of Schrödinger, from which he could derive very specific physical conclusions, namely the qualitative features of collision experiments (which were later verified quantitatively). This earned Born the Nobel prize in 1954 ”for his fundamental research in quantum mechanics, especially for his statistical interpretation of the wavefunction” [63].
Wallner’s Nobel prize presentation speech for the general public comments ([63]): ”Schrödinger had not solved the problem of how it is possible to make statements about the positions and velocities of particles if one knows the wave corresponding to the particle. Born provided the solution to the problem. He found that the waves determine the probability of the measuring results. For this reason, according to Born, quantum mechanics gives only a statistical description.” This is a bit inaccurate, as Born found probabilities for state transitions of particles, but not probabilities for positions and velocities of particles – this was done a little later by Pauli. This motivates a detailed study of the history of the statistical interpretation of quantum mechanics.
@@ 1.1 The Born rule before 1970
The Born rule has many slightly inequivalent formulations, ten different forms of which we state in the following, in modern terminology. Two formulations, the universal form (BR-US) and the discrete form (BR-DS) are taken almost verbatim from Wikipedia [79]. Four other formulations. are taken with minor changes from Chapter 14 of Neumaier [56] (but with the scattering form restated in two almost identical formulations with different meaning), and one more from Part II of Neumaier & Westra [58]. Two more general formulations were discovered only after 1970, and are described in Subsection 1.2. Finally, a tenth form (BR-Q) is presented in Subsection 1.3, adapted to quantum statistical mechanics and quantum field theory.
(BR-OSc), Born rule (objective scattering form): In a scattering experiment described by the S-matrix $S$,
@@
(BR-OE), Born rule (objective expectation form): The value of a quantity corresponding to an operator $X$ of a system in the pure state $\psi$ or the mixed state $\rho$ equals on average the quantum expectation value
@@
respctively. Here ${\mathop{\rm tr}\,}$ denotes the trace of an operator.
(BR-MSc), Born rule (measured scattering form): In a scattering experiment described by the S-matrix $S$, formula (1) gives the conditional probability density that when particles prepared in the in-state $\psi_{{{\mbox{\rm\scriptsize in}}}}$ are scattered, they are found in the out-state $\psi_{{{\mbox{\rm\scriptsize out}}}}$. Here the in- and out-states are asymptotic eigenstates of the Hamiltonian, labelled by a maximal collection of independent quantum numbers (including the energies, momenta and angular momenta of the particles involved).
@@
(BR-US), Born rule (universal spectral form): If a quantity corresponding to a self-adjoint Hermitian operator $X$ is measured in a system described by a pure state with normalized wave function $\psi$ then
@@
Since the formulation involves projection operators, measurements satisfying (BR-US) are called projective measurements; other name for these are von Neumann measurements and ideal measurements. Not all measurements are of this kind; see Subsection 1.2 for a more general class of measurements, and Subsection 3.4 for a fuller discussion of the practical limitations the class of projective measurements has.
(BR-FS), Born rule (finite spectral form): In a projective measurement of a Hermitian, self-adjoint operator $X$ (or vector of pairwise commuting operators) with a finite spectrum, the possible values are precisely the finitely many eigenvalues $\lambda_{k}$ of $X$ (or joint eigenvalues of the components of $X$), measured with a probability of
@@
(BR-DS), Born rule (discrete spectral form): Suppose that a quantity corresponding to a self-adjoint Hermitian operator $X$ with a discrete spectrum is measured in a system described by a pure state with normalized wave function $\psi$. Then:
@@
(BR-ME), Born rule (measured expectation form): If a quantity corresponding to a self-adjoint Hermitian operator $X$ is measured on a system in the pure state $\psi$ or the mixed state $\rho$, the results equal on average the quantum expectation value (2) or (3), respectively.
@@ 1.2 The Born rule after 1970
This straightforward generalization of the concept of a finite probability measure, given by a family of finitely many nonnegative numbers $p_{k}$ (probabilities) summing to 1, is a simplified version of a finite POVM. The following generalization of the finite spectral form (BR-FS) of the Born rule is essentially in Ali & Emch [2, (2.14)].
(BR-POVM), Born rule (POVM form): In a general quantum measurement, the possible values are finitely many distinct numbers or vectors $\lambda_{k}$, measured with a probability of (4), where the $P_{k}$ form a quantum measure.
@@
To justify (BR-POVM), the usual practice is to assume the Born rule for projective measurements as a basic premise, with a purely historical justification. Later (if at all), the more general quantum measure (POVM) setting is postulated and justified (e.g., in Fuchs & Peres [33]) in terms of the Born rule for projective measurements, in an extended state space formally constructed using an appropriate ancilla on the basis of Naimark’s theorem (Naimark [53]). This shows the consistency with the Born rule, but the ancilla has not always a meaning in terms of the physical Hilbert space.
Following Neumaier & Westra [58, Part II], we give in Section 3 precise definitions that lead to the correspondence between statistical expectations and quantum expectation claimed by the following condensed form of the Born rule.
(BR-C), Born rule (condensed form): The statistical expectation of the measurement results equals the quantum expectation of the measured quantity.
Both (BR-POVM) and (BR-C) are derived in Subsection 3.3 from a new, nontechnical measurement principle that is easy to motivate and understand, without any technicalities regarding the spectrum of operators.
@@ 1.3 Quantum expectation values
In quantum statistical mechanics and quantum field theory one usually makes a theoretical analysis of quantum processes where no actual measurement is made before the process is completed. Since one cannot invoke formulations of the Born rule that make reference to measurement, all probabilities and expectation values that appear in such an analysis, must be considered as formal expressions (2)–(4), with a clear mathematical meaning independent of the interpretation of quantum mechanics in terms of reality, knowledge, or measurement.
Every knowledge-based approach to statistical mechanics needs (BR-C) at the very basis, though this is usually left implicit and only tacitly assumed. Indeed, the only way to know something about quantum expectation values in an assumed model of a real quantum system is to gather information about the system or its source, and to translate it into information about the state of the system. The latter is usually done by means of the maximum entropy principle, which (Balian & Balazs [6]) directly equates the statistical expectation value with the quantum expectation value in order to be able to estimate the state. Note that even Heisenberg’s uncertainty relation is a statement about (theoretical) quantum expectation values that needs (BR-C) to imply anything about (measured) statistical uncertainty.
Since (BR-OSc) and (BR-OE) do not refer to measurement, they are not loaded with the difficulties surrounding the measurement problem of quantum mechanics. However, as we shall discuss in more detail in Subsection 2.3, were abandoned already by 1930, being considered as mathematically inconsistent. (A referee remarked that Bohmian mechanics restored the consistency of an objective, measurement-free view. But the restoration in Bohmian mechanics was only partial, namely for position, whereas there is still no objective interpretation of the measurement of energy and angular momentum. Thus (BR-OSc) and (BR-OE) do not hold.)
In particular, without objective interpretation, the quantum probabilities and quantum expectation values are purely theoretical quantities without any direct relation to reality. The link to reality is exclusively in the subject possessing the knowledge considered.
If one nevertheless wants to give the computations in quantum statistical mechanics an objective physical meaning, one must therefore define the objective quantum value of a vector $X$ of operators to be the quantum expectation value $\langle X\rangle$:
(BR-Q), Born rule (quantum value form): The quantum value $\overline{X}$ of a quantity corresponding to an operator $X$ of a system in the pure state $\psi$ or the mixed state $\rho$ is defined as the quantum expectation value (2) or (3), respectively. It is approximately measured, with an uncertainty of at least
@@
Here $X^{*}$ is the adjoint of $X$.
This form of the Born rule expresses the spirit of Heisenberg’s statement (in his 1927 paper on the uncertainty relation): ”To every quantum theoretical quantity or matrix, one may associate a number, that gives its ’value’, with a certain probable error; the probable error depends on the coordinate system; for every quantum theoretical quantity there is a coordinate system in which the probable error for this quantity vanishes.”
(German original, Heisenberg [35, p.181f]: ” Jeder quantentheoretischen Grösse oder Matrix lässt sich eine Zahl, die ihren ’Wert’ angibt, mit einem bestimmten wahrscheinlichen Fehler zuordnen; der wahrscheinliche Fehler hängt vom Koordinatensystem ab; für jede quantentheoretische Grösse gibt es je ein Koordinatensystem, in dem der wahrscheinliche Fehler für diese Grösse verschwindet.”) (Heisenberg [35, p.181f])
@@
In quantum field theory, quantum expectation values of non-Hermitian operators are routinely used; they are already needed in the definition of so-called $N$-point functions as quantum expectation values of products of field operators. That $2$-point functions are in principle observables through linear response theory shows that the operators corresponding to observable quantities need not be Hermitian. Indeed, Dirac’s 1930 quantum mechanics textbook (Dirac [24, p.24]), which introduced the name ”observable” for operators, used this terminology for arbitrary linear operators: ”It is convenient to count any operator that can be multiplied into the $\psi$s and $\phi$s in accordance with the foregoing axioms as an observable.”
It is therefore interesting to note that all later editions of Dirac’s textbook and all later textbooks on quantum mechanics require the restriction to Hermitian operators possessing a real spectral resolution, i.e., in modern terminology, to self-adjoint Hermitian operators. Probably Dirac became aware of the fact that when $X$ is not normal (and in particular when $X$ is defective, hence has not even a spectral resolution), it is impossible to give the quantum value $\langle X\rangle$ of $X$ a spectral interpretation – which before 1970 was an essential ingredient of all measurement-based forms of the Born rule. Only the more recent formulations (BR-POVM), (BR-C), and (BR-Q) are again free of a spectral restriction.
@@ 1.4 Discussion
tacitly made in classical statistical physics, and imported from there into the spectral forms of the Born rule. In practice, we only measure a small set of distinguished variables suc has distances, position, energy, momentum, angular momentum, spin, helicity, and not arbitrary nonlinear expressions in these. Hence it is quite reasonable to require (D) not for arbitrary nonlinear $f$ but only for linear $f$.
Note that giving up (D) is necessary if one wants to account for experiments involving the simultaneous measurement of position and momentum, such as that routinely used in modern collision experiments, where particle tracks with fairly well-determined positions and momenta are reconstructed in in a time projection chamber. A POVM representation of the latter is given in Neumaier & Westra [58, Section 6.2.4]. Indeed, (D) is also incompatible with (BR-POVM).
The three newest forms of the Born rule, (BR-POVM), (BR-C) , and (BR-Q), are mathematically much more elementary since they do not depend on the notion of self-adjointness of operators. The POVM form (BR-POVM) of the Born rule is more or less equivalent to the condensed form (BR-C); see Subsection 3.3. As already stated in Subsection 1.2, (BR-POVM) is a generalization of the finite spectral form (BR-FS).
@@
von Neumann [60, p.255] derives for a quantum expectation value satisfying four (for von Neumann) plausible conditions (conditions A.–D. listed and motivated on p.250–252) the necessity of the formula (3) with a Hermitian density operator $\rho$ (his $U$) of trace 1. This is abstract mathematical reasoning, still within the core of formal (uninterpreted) quantum mechanics, independent of any relation to measurement. Note that assuming von Neumann’s condition D. which is a restricted form of condition (D) above, forces the measurements to be projective measurements. Thus he misses the POVM generalization of the Born rule since the latter violates his condition D.